OpenStudy (anonymous):

(2a+4)(4a+7), I can get the first part, 8a^2, and the last part, 28, but i have forgotten how to get the middle part...

OpenStudy (anonymous):

FOIL - first, outer, inner, last. multiplying the first two terms together 2a*4a=8a^2, multiply the two outer terms next, 2a*7 = 14a, then the two inner terms, 4*4a = 16a, and the last terms: 4*7 = 28, so you get (2a+4)(4a+7) = 8a^2 + 14a + 16a + 28 = 8a^2 + 30a + 28

OpenStudy (anonymous):

I remeber how to do foil, but I do remember how to do the box method

OpenStudy (anonymous):

that anser was wrong. it was 8a^2 +30a+28

OpenStudy (anonymous):

that's what i said?